OpenStudy (anonymous):

If p is directly proportional to q and if p = 14 when q = 8, what is the value of p when q = 28

OpenStudy (anonymous):

p=34 although that may not be right

OpenStudy (texaschic101):

14/8 = p/28 cross multiply (8)(p) = (28)(14) 8p = 392 p = 392/8 p = 49 check.. 14/8 = 49/28 cross multiply (8)(49) = (28)(14) 392 = 392 (correct) p = 49 when q = 28

OpenStudy (anonymous):

you may do like this also. If p is directly proportional to q, then p=k q If p is inversely proportional to q,then p=q/k where k is a constant. 14=8k,k=14/8=7/4 \[p=\frac{ 7 }{4 }q,when q=28,p=\frac{ 7 }{4}*28=49\]